9.74. Suppose we toss a biased coin independently until we get two heads or two tails in total. The coin produces a head with probability p on any toss.

1. What is the sample space of this experiment?
2. What is the probability function?
3. What is the probability that the experiment stops with two heads?
Answer #1

9.74)

Now, P(head) = p and P(tail) = 1-p

1)

The Sample space is given by:

S=\begin{Bmatrix} HH,TT,HTH,THH,THT,HTT \end{Bmatrix}

2)

P(HH)=p.p=p^2,P(TT)=(1-p)(1-p)=(1-p)^2

P(HTH)=p(1-p)p=p^2(1-p),P(THH)=(1-p)p.p=p^2(1-p)

P(THT)=(1-p)p(1-p)=p(1-p)^2,P(HTT)=p(1-p)(1-p)=p(1-p)^2

So, the probability function is given by:

X P(X)
HH p^2
TT (1-p)^2
HTH p^2(1-p)
THH p^2(1-p)
THT p(1-p)^2
HTT p(1-p)^2

3)

Required probability = P(experiment stops with two heads)

= P(HH) + P(HTH) + P(THH)

p^2+p^2(1-p)+p^2(1-p)=p^2(3-2p)